OpenStudy (anonymous):

When does a fraction have a value of 1?

OpenStudy (studygurl14):

When the numerator and the denominator are the same number. For example: 4/4 = 1 5/5 = 1 2/2 = 1

OpenStudy (anonymous):

when the top and bottom are the same